# Structured Packing in LLM Training Improves Long Context Utilization

### v6

@@ Appendix B Detailed Results For Medium-Size Models
The main advantage of SPLiCeBM25/SPLiCeCont over the SPLiCeRepo approach is that it can be used also for non-structured data. Table 12 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Table 13 considers models trained with YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023) and Naive (no adjustment to RoPE) context extension methods. Table 14 shows that a simple artificial extension of example length via random concatenation of documents does not help.
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways. For details about training please refer to Section 3.1.
@@
Table 14: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways. For details about training please refer to Section 3.1.
@@
Eval/Method SPLiCeBM25 Baseline Random
@@ C.1 SPLiCe Parameters
C.2 Importance of Separating Tokens
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training examples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement). The results are shown in Table 16. We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains.
Table 16: Perplexity evaluation of two methods of organizing the data. Baseline – document equals training example. Random – concatenate documents into examples of length bounded by $120$K characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in Random does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation.
Long Context Method arXiv Code Code &
@@ Appendix D Key-Value Retrieval Task
Figure 2 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in (Liu et al., 2023). This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
@@
Figure 3: Performance on a smaller version of key-value retrieval task from (Liu et al., 2023). We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama (Rozière et al., 2023) method of context extension. Each position was evaluated using $500$ examples.
@@ Appendix E Perplexity Improvements
In Figure 4 we present perplexity improvements of $3$B FoT SPLiCeBM25 over Baseline. Figure 5 shows the evolution of SPLiCe model perplexity during the training. We follow (Anthropic, 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
@@ Appendix F Shuffling
Table 17: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001). We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was an identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
@@ G.1 Evaluation Data
Table 18: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents having not less than $c$ tokens and extract the $c$ tokens prefix.
@@ Appendix H Faiss Parameters
Our experiments with SPLiCeCont utilize Faiss (Johnson et al., 2019) for fast approximate inner-product search. To be more precise, we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
@@ Appendix I Detailed Accuracy Improvements
Figure 6 shows details about accuracy improvements on TREC when considering different numbers of in-context examples.
@@
In Table 19, we present the perplexity results for $270$M parameter models trained as described in Section 3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
@@
Long Context Method arXiv Code Code &
@@
Chan et al. (2022); Han et al. (2023) shows that the ”burstiness” property of training data benefits a model’s generalisation. Specifically, ”burstiness” presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency. We also observe a burstiness property of data in SPLiCe, as shown in Table 20.

### v7

@@ B.3 Detailed results
The main advantage of SPLiCe BM25 /SPLiCe Cont over the SPLiCe Repo approach is that it can also be used for non-structured data. Table 12 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Tables 13, 14 consider models trained with YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023] and Naive (no adjustment to RoPE) context extension methods. Table 15 shows that a simple artificial extension of example length via random concatenation of documents (C source files) does not help.
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways. For details about training, please refer to Section 3.3.
@@
Table 14: Perplexity results comparing different ways of organizing the same data for non-FoT models. All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama with a Naive context extension method.
Altered Train Data: C#
@@
Table 15: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways.
@@
Eval/Method SPLiCe BM25 Baseline DomRnd
@@ C.1 SPLiCe parameters
C.2 Importance of separating tokens and in-domain sampling
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training examples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens and concatenation of documents from the same domain, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the C dataset (sampling without replacement). The results are shown in Table 17. We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains. When concatenating the documents, we concatenate them within the domain (C language in this case).
Table 17: Perplexity evaluation of two methods of organizing the data. Baseline – example packing, document equals training example. DomRnd – concatenate documents within the domain into examples of length bounded by $120$K characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in DomRnd does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation. The training data is a 50/50 mixture of RedPajama, organized in a standard way, and C data, organized using a method of choice.
Altered Method arXiv Code Code &
@@ Appendix D Key-Value retrieval task
Figure 2 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in [Liu et al., 2023]. This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
@@
Figure 3: Performance on a smaller version of key-value retrieval task from [Liu et al., 2023]. We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama [Rozière et al., 2023] method of context extension. Each position was evaluated using $500$ examples.
@@ Appendix E Perplexity improvements
In Figure 4 we present perplexity improvements of $3$B FoT SPLiCe BM25 over Baseline. Figure 5 shows the evolution of SPLiCe model perplexity during the training. We follow [Anthropic, 2023] and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
@@ Appendix F Shuffling
Table 18: Average classification performance on TREC [Li and Roth, 2002, Hovy et al., 2001]. We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was the identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
@@ G.1 Evaluation data
Table 19: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents that have not less than $c$ tokens and extract the $c$ tokens prefix.
@@ Appendix H Faiss parameters
Our experiments with SPLiCe Cont utilize Faiss [Johnson et al., 2019] for fast approximate inner-product search. To be more precise, we use the "IVF8192,Flat" index that we train on 262144 examples coming from the dataset.
@@ Appendix I Detailed accuracy improvements
Figure 6 shows additional details about accuracy improvements on TREC when considering different numbers of in-context examples.
@@
In Table 20, we present the perplexity results for $270$M parameter models trained as described in Section 3.3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
@@
Altered Method arXiv Code Code &
@@
Chan et al. [2022], Han et al. [2023] shows that the "burstiness" property of training data benefits a model’s generalisation. Specifically, "burstiness" presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency. We also observe a burstiness property of data in SPLiCe, as shown in Table 21.
@@
Appendix L HumanEval
Table 22: We additionally evaluate our $7$B and $13$B parameter models on HumanEval [Chen et al., 2021]. For $7$B models, all methods experience a decrease in performance, but SPLiCe is better at maintaining short context performance. For $13$B, both SPLiCe and Baseline increase the performance. We note that in our evaluation we do not use special tokens designed for CodeLlama $13$B code completion. Instead, evaluate all models in the same way. Because of the different evaluation pipeline our results for Starting Checkpoint are not the same as the ones presented in [Rozière et al., 2023].
Model Method Pass@1
$7$B Starting Checkpoint 12.9
[…]
Baseline 9.1
[…]
Baseline 8.5
$13$B Starting Checkpoint 23.8
[…]
Baseline 26.8